OpenStudy (anonymous):

what is the sum of these mixed numbers ? 3 1/2 + 6 1/7 how would i find this anwer my options are 9 5/7 9 4/7 9 9/14

OpenStudy (amistre64):

since 6+3 = 9; that part is common to them all so id ignore it and see what 1/2+1/7 turns out to be

OpenStudy (anonymous):

It would 9 9/14. 1/2+1/7, the common denominator would be 14. ill let you figure out the rest ;)

OpenStudy (anonymous):

3 1/2 + 6 1/7 = 135/14 = 9.642857142857142